I need help with this problem:
x+y=500 and 5x+10y=3000.
I need to use elimination to solve this problem.
I would really appreciate the help(:

Answer by Maths68(1474)

x+y=500...........(1)
5x+10y=3000.......(2)

Take (2)
5(x+2y)=5(600)
Divide by 5 both sides
5(x+2y)/5=5(600)/5
x+2y=600...............(3)

subtract (1) from (3)
x+2y=600...............(3)
-x-y=-500..............(1)
--------------------------
y=100
Put the value of y in (1)
x+y=500...........(1)
x+100=500
x=500-100
x=400
x,y=400,100
Check
=====
Put the value of x and y (1)
x+y=500...........(1)
400+100=500
500=500
Put the value of y in (2)
5x+10y=3000.......(2)
5(400)+10(100)=3000
2000+1000=3000
3000=3000